Hi, I dont know what Im doing wrong I need help please. Here is the question: One year ago you invested $4300 into three accounts that earn 3.25%, 4%, and 5.75% annual interest. The interest you earned after one year is $193. You remembered you invested three times as much in the 4% account than you did the 3.25% account. How much did you invest in each type of account? This is what i have so far, I dont think its right.
x + y + Z = 4300
325x + 400Y + 575z = 19300
Y=3x
x + (3x) + Z = 4300
4x + z = 4300
325x + 400 (3x) + z =19300
1525x +575z = 19300
(-575) 4x + z = 4300 (-575)
-2300x - 575z =-2472500
1525x + 575z = 19300
-775x = -2453200
x=3165.4 then when I do 3 times x to get y it doesnt fit with the total of 4300. I dont know please help me thank you.

Answer by stanbon(57347)

Hi, I dont know what Im doing wrong I need help please. Here is the question: One year ago you invested $4300 into three accounts that earn 3.25%, 4%, and 5.75% annual interest. The interest you earned after one year is $193. You remembered you invested three times as much in the 4% account than you did the 3.25% account. How much did you invest in each type of account? This is what i have so far, I dont think its right.
x + y + Z = 4300
0.0325x + 0.0400Y + 0.0575z = 193
Y=3x
-----------
x + y + z = 4300
325x + 400y + 575z = 1930000
3x - y + 0 = 0
------------
Solve the system of equations by any method to get:
x = 700
y = 2100
z = 1500
===========
Cheers,
Stan H.